OpenStudy (anonymous):

A QUESTION FOR THE MASTERS OF MATH A MEDAL WILL BE REWARDED AND I WILL FAN YOU…ARE YOU UP FOR THE CHALLENGE: 2x-9/x^2- x- 6 = A/x-3 + B/x+2

OpenStudy (anonymous):

MWAHAHAHA

OpenStudy (anonymous):

@ganeshie8 @midhun.madhu1987 @kropot72

OpenStudy (anonymous):

YOU HAVE BEEN CHOSEN! (DUN DUN DUUUUUN)

OpenStudy (anonymous):

I HAVE A DETAILED EXPLANATION AND ANSWERS DO YOU?

OpenStudy (anonymous):

Does the equation look like that?\[\frac{ 2x-9 }{ x^2 -x-6 }=\frac{ A }{x-3 }+\frac{ B }{ x+2 }\]

OpenStudy (anonymous):

@ganeshie8 @midhun.madhu1987 @kropot72

OpenStudy (midhun.madhu1987):

is the question the same as drawn by @faith2012 ?

OpenStudy (anonymous):

@faith2012 HAS THE EQUATION CORRECT

OpenStudy (anonymous):

I SWEAR I HAVE THE ANSWER THIS IS TO PUSH YOUR SKILL TO THE LIMIT YOU WILL NOT BE REPORTED FOR GIVING THE FULL ANSWER

OpenStudy (anonymous):

TELL ME WHEN YOU GIVE UP AND I WILL REVEAL THE ANSWER

OpenStudy (anonymous):

MWHAHAHAHA

OpenStudy (midhun.madhu1987):

 try to split them and u get the Right side